# LPSim: Large Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework

### v1

@@ Abstract
Traffic simulation is a tool for congestion analysis, travel time estimation, and route optimization for urban planning. It is beneficial for navigation apps, transportation network companies, and state agencies. Traditionally, the traffic micro-simulation frameworks are based on road segments and can only support several main roads. Efficient traffic simulation on a regional scale remains a significant challenge due to the complexity of urban mobility and the large scale of spatiotemporal data. This paper introduces a Large Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework (LPSim), which is a novel multi-modal regional scale traffic simulation framework that leverages large-scale graphical processing unit (GPU) parallel computing to address these challenges. LPSim utilizes a multi-GPU architecture to simulate extensive and dynamic traffic networks with high fidelity and reduced computation time. Using the parallel processing capabilities of GPUs, LPSim can perform tens of millions of individual vehicle dynamics simulations simultaneously, significantly outperforming traditional CPU-based approaches. The framework is designed to be scalable and can easily accommodate the increasing complexity of traffic simulations. We present the theory behind GPU-based traffic simulation, the architecture of single- and multiple-GPU-based simulations, and the graph partition strategies that optimize computation resource allocation. Our experimental results demonstrate the effectiveness of LPSim in simulating large-scale traffic scenarios. LPSim is capable of completing simulations of 2.82 million trips in merely 6.28 minutes on a single GPU machine equipped with 5120 CUDA cores (Tesla V100-SXM2). Furthermore, utilizing an AWS p2 instance with dual NVIDIA K80 GPUs, which collectively offer 4992 CUDA cores, LPSim successfully simulates 9.01 million trips within 21.45 minutes. This performance not only demonstrates its speed and scalability advantages over traditional simulation techniques but also highlights LPSim’s unique position as the first traffic simulation framework that is adaptable across multi-modal (multiple transportation modes e.g., bikes, subways, and buses) contexts and scalable for both single and multiple GPU configurations. Consequently, LPSim provides an invaluable tool for individuals and extensive research teams alike, enabling the acquisition of large-scale traffic simulation results in a time-efficient manner. LPSim code is available at: https://github.com/Xuan-1998/LPSim
@@
Regional-scale traffic simulation framework, GPU Parallel Computing, Graph Partitioning, Roofline Model
@@ 1 Introduction
Traffic simulation, a powerful tool that bridges the areas of computer science and traffic engineering, plays a crucial role in understanding how urban mobility works Zomer et al. [2015]Jiang et al. [2023]. By simulating and recreating traffic scenarios in virtual environments, traffic simulation systems offer a robust platform for researchers, engineers, and policymakers to analyze, experiment, and design strategies to avoid the risks and costs of real-world trials Krautter et al. [1999].
@@
One prevalent approach in traffic simulation is the use of microscopic simulation Treiber et al. [2000] Maroto et al. [2006], which refers to a computer-based modeling technique that simulates the behavior and interactions of individual entities at a microscopic level. In the traffic field, microscopic simulation takes a granular perspective, focusing on individual vehicle movements, accelerations, and lane changes to provide a realistic representation of traffic flow. Compared to macroscopic simulation and mesoscopic simulation, which focus on complete road flow instead of individual vehicle movement Helbing and Molnar [1995], people often use microscopic simulation on the level of cities and highways Maroto et al. [2006] Du et al. [2015], which are smaller scale. However, in recent years, the popularity of large-scale microscopic simulation has grown significantly, driven by the alignment of time-driven microsimulation with the single instruction, multiple data (SIMD) nature of GPU Yedavalli et al. [2021], and the increasing demand for techniques to accommodate various new traffic modes. For example, the Urban Air Mobility (UAM) system Muna et al. [2021] has a regional impact, requiring researchers to evaluate its performance.
@@
Compared with regular-scale microscopic simulation, the aforementioned large-scale microscopic traffic simulation has its challenges. In particular, modeling individual vehicle movements in fine detail, continuously processing, updating the states of numerous individual vehicles, and managing extensive spatiotemporal data in the large-scale microscopic simulation will lead to high computational requirementsAlgers et al. [1997]. Therefore, for practitioners and researchers, large-scale, time-varying simulation data storage and update call for efficient computation tools and a scalable storage approach.
@@
In traffic simulation, the development of GPUs makes it possible to manage the intricate dynamics of vehicle movements and large-scale data structures. Thus, it is possible to utilize the developed GPU architectures to enhance traffic simulation Yedavalli et al. [2022]. In the new traffic simulation framework, LPSim, presented in this paper, we use GPUs to handle both the spatial and temporal aspects of large-scale microscopic traffic simulations. Spatial aspects include the addition or deletion of vehicles, the transfer of vehicle statuses (start, on route, stop) of independent vehicles, and the spatial propagation of the vehicle within the network Barceló et al. [2010]Boxill and Yu [2000]. The temporal dynamics mainly consist of the time propagation and synchronization of traffic flow for each time stepOsorio and Nanduri [2015]. However, the challenge of GPU memory limitations, with a single GPU’s memory capped at 80 GB (for A100) Choquette and Gandhi [2020] and most GPUs have only 8GB memory Zhang et al. [2017], contrasts with the potential for Central Processing Units (CPUs) to scale up to 768 GB (for AWS) Pelle et al. [2019]. To overcome this, our framework employs graph partitioning methods to distribute vast amounts of transportation network and vehicle movement data across multiple GPUs. This strategy ensures that the simulation can scale to accommodate large-scale networks without compromising the level of detail or simulation speed. Our main design innovation for the single GPU framework is summarized in Table 1, and for the multi-GPU system is summarized in Table 2. Based on our research, we claim the contributions in the following four aspects:
@@
    Vectorized Data Storage and Access Implementation: Our approach incorporates vectorized data storage and access mechanisms that allow for the efficient storage and access of both transportation network data and vehicular movement/speed information within a GPU environment. This innovation facilitates improved data handling and processing speed within the GPU architecture.
@@
    Design of the Multi-GPU Transportation Simulation Framework: We have developed a transportation simulation framework that uses multiple GPUs, structured around the concept of graph partitioning. This design ensures the results don’t change as the number of GPUs increases, thereby enhancing the effectiveness and efficiency of the simulation.
@@
    Scalability Benchmarking: The scalability of LPSim has been thoroughly evaluated both theoretically and through empirical research on both one GPU and multiple GPUs. Scalability assessments were conducted using data sourced from the San Francisco County Transportation Authority (SFCTA), complemented by performance analysis based on the Bulk Synchronous Parallel Model Gerbessiotis and Valiant [1994], Amdahl’s law Hill and Marty [2008], and Roofline ModelWilliams et al. [2009], . LPSim demonstrates remarkable efficiency, completing simulations of 2.82 million trips in just 6.28 minutes on a single GPU setup featuring a Tesla V100-SXM2 with 5120 CUDA cores compared with the 12 hours computation time for 0.6 million trips in the same area (San Francisco Bay Area) using CPU-based(conventional) simulation Hsueh et al. [2021], who is simulating less demand but takes longer simulation time and lower simulation fidelity with a mesoscopic simulation approach. Additionally, when the scale becomes larger and has to be deployed on an AWS p2 instance equipped with dual NVIDIA K80 GPUs, which together provide 4992 CUDA cores, LPSim can efficiently simulate 9.01 million trips in a mere 21.45 minutes. These analyses demonstrate the framework’s ability to scale effectively in response to varying computational demands.
  • 4.
    Trade-off between Multi-GPU Parallel Computing and Communication Time: Our investigations reveal that while multi-GPU parallel computing can expedite the simulation process, it is also subject to potential slowdowns due to increased communication times among GPUs. Experiments conducted on systems with 2, 4, and 8 GPUs on a gcloud instance with NVIDIA V100 GPUs have yielded the data, as detailed in Table 3, which is explained further in Section 4.2. It indicates that the increase in the number of GPUs, coupled with appropriate graph partitioning strategies, can lead to a reduction in total computation time required for simulation.
@@
Traffic Data Structure Architecture Design
Network Large scale datab Contiguous memory blocks Fig 2
@@
Agent status Eq. 1 Markov Update GPU SIMD Yilmazer et al. [2014]
  • b
    The network contains more than half a million edges.
@@
Multi GPU challenge Architecture Design
Multi GPU load balancing Graph partitioning Fig 4
Inter GPU communication Ghost zone design Fig 3
Multi GPU synchronization GPU device based vector Fig 8
@@
Table 3: Simulation Time /ms
GPU Numbers 2 4 8
Random Partition Dumped Dumped Dumped
Balanced Partition 2498466 1483472 1269893
Unbalance Partition 2014554 1679861 1783668
@@ 3.1 Basic Components
$l_{\text{min}}$ Minimum length of the link of the road network $v_{i}$ Speed of mode (car, bike, rail, etc.) $i$
@@
$u_{i}(k)$ The velocity of vehicle $i$ at time step $k$ $f_{i}$ The function abstraction of IDM car following, lane change, and gap acceptance
$m_{i}$ Probability of a mandatory lane change for vehicle $i$ $x_{i}$ Distance of vehicle $i$ to an exit or intersection
$x_{0}$ Distance of a critical location to the exit or intersection $g_{\text{lead}}$ Critical lead gap for a lane change
$g_{\text{lag}}$ Critical lag gap for a lane change $g_{a}$ Desired lead gap for a lane change
@@
$T$ Desired time headway
In this part, we introduce the vehicle dynamics used for microsimulation initially developed in Yedavalli et al. [2021]. There are three key components: intelligent driver model, lane change process, and gap acceptance.
We further introduce the concept of Time Step Criteria (TSC) to enhance the model’s adaptability to a multimodal traffic simulation framework. Additionally, we present a theory of parallelizable simulation that aligns with the GPU’s SIMD architecture as shown in Eq. 1. This includes a detailed explanation of how memory is utilized not only to represent the road network but also to depict the occupancy and speed of the vehicles. Lastly, we consolidate our methodology into a comprehensive algorithm, outlined as Algorithm 1. Readers can refer to table 4 for notation in this part.
First, we introduce a set of related notations in Table 4. The dynamics of vehicles within our simulation Treiber and Kesting [2013], is governed by the Intelligent Driver Model (IDM) Albeaik et al. [2022], which models the acceleration of vehicles for updating vehicle locations at each time step. Second, the mandatory lane change process is encapsulated by the formula (Lane Change), as described in Iqbal et al. [2014]. For the vehicle, when the distance to an exit falls below a threshold distance $x_{0}$, it is triggered to make a mandatory lane change and the probability of such a change increases as the vehicle approaches the exit. Once a vehicle inquires for a lane change, it must assess the feasibility of this maneuver by assessing the gaps with both the leading and the lagging vehicles. This assessment is carried out according to the formula (Gap Acceptance), following the guidelines set forth by Choudhury et al. [2007].
$\dot{v}=a\left(1-\left(\frac{v}{v_{o}}\right)^{\delta}-\left(\frac{s_{o}+Tv+\frac{v\Delta v}{2\sqrt{ab}}}{s}\right)^{2}\right)$ (IDM)
$m_{i}=\begin{cases}e^{-\left(x_{i}-x_{0}\right)^{2}}&x_{i}\geq x_{0}\\ 1&x_{i}\textless x_{0}\end{cases}$ (Lane Change)
$\begin{cases}\par g_{lead}&=\max\left(g_{a},g_{a}+\alpha_{a}v_{i}+\alpha_{i}\left(v_{i}-v_{a}\right)\right)+\epsilon_{a}\\ g_{lag}&=\max\left(g_{b},g_{b}+\alpha_{i}v_{i}+\alpha_{b}\left(v_{i}-v_{b}\right)\right)+\epsilon_{b}\end{cases}$ (Gap Acceptance)
Finally, the time step criteria (TSC) means that the relationship between the time step and the spatial resolution needs to be constrained in the following way. For any time step $i$:
$v_{i}\times T_{tsc}\geq S_{tsc}$ (TSC-1)
$V_{i}\times T_{tsc}\leq l_{\text{min}}$ (TSC-2)
Inequality TSC-1 guarantees that each vehicle advances by at least one cell in every time step. Conversely, inequality (TSC-2) ensures that the distance covered by any vehicle in a single time step does not exceed the length of the shortest link in the road network.
@@
$l_{i}(k+1)=f_{i}(\underbrace{l_{i}(k),u_{i}(k)}_{\text{Current Vehicle}},\underbrace{l_{i-n}(k),\ldots,l_{i+n}(k)}_{\text{Surrounding Vehicles}},\underbrace{u_{i+1}(k)}_{\text{Front Vehicle}})$ (1)
@@
Remark: One byte can only be occupied by one vehicle, and we are not simulating the case that vehicles crash with each other.
As shown in Figure 2, the whole network will be represented in GPU memory with Char type in C++ which can represent Numbers from 0 to 255, we use it in the following meaning:
@@
    1 Byte in Memory = 1 Meter in Road Segments
@@
    Value 255 in Memory = Not Occupied by anything in Road Segments
@@
    Value 0-254 = Occupied by Vehicle and the Value is the Speed of the Vehicle in m/s
Figure 2: One GPU representation

### v2

# Large Scale Multi-GPU Based Parallel Traffic Simulation for Accelerated Traffic Assignment and Propagation

@@ Abstract
Traffic simulation is a critical tool for congestion analysis, travel time estimation, and route optimization in urban planning, benefiting navigation apps, transportation network companies, and state agencies. Traditionally, traffic micro-simulation frameworks are based on road segments and can only support a limited number of main roads. Efficient traffic simulation on a regional scale remains a significant challenge due to the complexity of urban mobility and the large scale of spatiotemporal data. This paper introduces a Large Scale Multi-GPU Parallel Computing based Regional Scale Traffic Simulation Framework (LPSim), which leverages graphical processing unit (GPU) parallel computing to address these challenges. LPSim utilizes a multi-GPU architecture to simulate extensive and dynamic traffic networks with high fidelity and reduced computation time. Using the parallel processing capabilities of GPUs, LPSim can perform tens of millions of individual vehicle dynamics simulations simultaneously, significantly outperforming traditional CPU-based approaches. The framework is designed to be scalable and can easily accommodate the increasing complexity of traffic simulations. We present the theory behind GPU-based traffic simulation, the architecture of single- and multi-GPU based simulations, and the graph partition strategies that enhance computation resource load balance. Our experimental results demonstrate the effectiveness of LPSim in simulating large-scale traffic scenarios. LPSim is capable of completing simulations of 2.82 million trips in just 6.28 minutes on a single GPU machine equipped with 5120 CUDA cores (Tesla V100-SXM2). Furthermore, utilizing a Google Cloud instance with two NVIDIA V100 GPUs, which collectively offer 10240 CUDA cores, LPSim successfully simulates 9.01 million trips within 21.16 minutes. We further tested our simulator with the same demand on dual NVIDIA A100-PCIE-40GB GPUs, which finished the simulation in 0.0398 hours, approximately 113 times faster than the same simulation scenario running on an Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz, which takes 4.49 hours to complete. This performance not only demonstrates its speed and scalability advantages over traditional simulation techniques but also highlights LPSim’s unique position as the first traffic simulation framework that is scalable for both single- and multiple-GPU configurations. Consequently, LPSim provides an invaluable tool for individuals and extensive research teams alike, enabling the acquisition of large-scale traffic simulation results in a time-efficient manner. LPSim code is available at: https://github.com/Xuan-1998/LPSim
@@
Regional-scale traffic simulation framework, GPU Parallel Computing, Graph Partitioning, Roofline Model, Digital Virtual Transport
@@ 1 Introduction
Traffic simulation, a powerful tool that bridges the areas of computer science and traffic engineering, plays a crucial role in understanding how urban mobility works Zomer et al. [2015]Jiang et al. [2023]. By simulating and recreating traffic scenarios in virtual environments, traffic simulation systems offer a robust platform for researchers, engineers, and policymakers to analyze, design, and experiment strategies to avoid the risks and costs of real-world trials Krautter et al. [1999].
@@
A prevalent approach in traffic simulation is the use of microscopic simulation Treiber et al. [2000] Maroto et al. [2006], which refers to a computer-based modeling technique that simulates the behavior and interactions of individual entities at a microscopic level, which models detailed physical dynamics between vehicles. In the traffic field, microscopic simulation takes a granular perspective, focusing on individual vehicle movements, accelerations, and lane changes to provide a realistic representation of traffic flow. Compared to macroscopic simulation and mesoscopic simulation, which focus on complete road flow instead of individual vehicle movement Helbing and Molnar [1995], people often use microscopic simulation on the level of cities and highways Maroto et al. [2006] Du et al. [2015], which are smaller scale. However, in recent years, the popularity of tens of millions of trips of large-scale microscopic simulation has grown significantly Maciejewski et al. [2016], and the increasing demand for techniques to accommodate various new traffic modes. For example, the Urban Air Mobility (UAM) system Muna et al. [2021] has a regional impact, requiring researchers to evaluate it at a regional scale. Yedavalli et al. [2021] tried to do regional scale simulation with single GPU and can scale up to 3.2M trips. With the same network and demand used by Yedavalli et al. [2021], Figure 1 shows that in the same simulation scale in SF bay area with 3.2M OD trips LPSim stands out to be the fastest simulator of all the Traffic Simulators. Other than the simulation speed, LPSim can scale the simulation up to 24M OD trips and beyond as shown in Figure 13, which none of the simulators can achieve yet.
Figure 1: Simulation Time Comparison for Regional Traffic Simulation
@@
Compared to regular-scale microscopic simulation, the aforementioned large-scale microscopic traffic simulation has its challenges. In particular, modeling individual vehicle movements in fine detail, continuously processing, updating the states of numerous individual vehicles, and managing extensive spatiotemporal data in the large-scale microscopic simulation will lead to high computational requirementsAlgers et al. [1997]. Therefore, for practitioners and researchers, large-scale, time-varying car following model based traffic assignment and propagation call for efficient computation tools and scalable framework.
@@
It is possible to utilize the developed GPU architectures to improve traffic simulation Yedavalli et al. [2022]. LPSim, presented in this paper, uses GPUs to handle both the spatial and temporal aspects of large-scale microscopic traffic simulations. Spatial aspects include the network partitioning into different GPUs, GPU threads Reallocation between GPUs, and Ghost Zone Creation between GPUsBarceló et al. [2010]Boxill and Yu [2000]. The temporal dynamics mainly consist of the time-driven propagation and synchronization across time steps of traffic flow for each time stepOsorio and Nanduri [2015]. However, the challenge of GPU memory limitations, with a single GPU’s memory capped at 80 GB (for A100) Choquette and Gandhi [2020] and most GPUs have only 8GB memory Zhang et al. [2017], contrasts with the potential for Central Processing Units (CPUs) to scale up to 768 GB (for AWS) Pelle et al. [2019]. To overcome this, our framework employs graph partitioning methods to distribute vast amounts of transportation network and vehicle movement data across multiple GPUs. This strategy ensures that the simulation can scale to accommodate large-scale networks without compromising the level of detail or simulation speed. And we used the combination of car following model, lane change model, and gap acceptance model to model traffic propagation which is a common approach used by popular micro simulators, such as SUMO Krajzewicz et al. [2002], Vissim Lownes and Machemehl [2006], Aimsun Casas et al. [2010], we used Intelligent Driver Model (IDM) Kesting et al. [2010] and our framework is adaptable to any other types of traffic propagation models. We tested our simulator on an Intel(R) Xeon(R) Gold 6326 CPU @ 2.90GHz with a typical day demand of 9008766 trips from 0 AM to 12 PM provided by SFCTA Outwater and Charlton [2006] on a SF bay area road network with 224,223 nodes and 549,008 edges which finished with 4.49192555556 hours compared to a dual NVIDIA A100-PCIE-40GB GPUs’ 0.0398483333 hours simulation time which are around 113 times faster. Our main design innovation for the single GPU framework is summarized in Table 1, and for the multi-GPU system is summarized in Table 2. Based on our research, we claim contributions in the following three computational aspects, rather than theoretical transportation models:
@@
    Design of the Multi-GPU Transportation Simulation Framework: We have developed a transportation simulation framework that uses multiple GPUs, structured around the concept of graph partitioning. This design ensures the simulation outputs remain consistent as the number of GPUs increases, thereby enhancing the effectiveness and efficiency of the simulation.
    • (a)
      Vectorized Data Storage and Access Implementation: Our approach incorporates vectorized data storage and access mechanisms that allow for the efficient storage and access of both transportation network data and vehicular movement/speed information within a GPU environment. This innovation facilitates improved data handling and processing speed within the GPU architecture, which are detailed explained in Figure 4.1.
@@
    Scalability Benchmarking: The scalability of LPSim has been thoroughly evaluated both theoretically and through empirical research on multiple GPUs with Google Cloud Service and generic personal GPUs. Scalability assessments were conducted using data sourced from the San Francisco County Transportation Authority (SFCTA), complemented by performance analysis based on the Bulk Synchronous Parallel Model Gerbessiotis and Valiant [1994], Amdahl’s law Hill and Marty [2008], and Roofline Model Williams et al. [2009], LPSim demonstrates remarkable efficiency, when the scale becomes larger and has to be deployed on an Google Cloud instance equipped with V100 GPUs, LPSim can efficiently simulate 9.01 million trips in 21.16 minutes, and gains nearly 1.68 times speedup when increases the number of GPUs from 2 to 4 with balanced partition of the road network. These analyses demonstrate the framework’s ability to scale effectively in response to varying computational demands.
@@
    Trade-off between Multi-GPU Parallel Computing and Communication Time: Our investigations reveal that while multi-GPU parallel computing can expedite the simulation process, it is also subject to potential slowdowns due to increased communication times among GPUs. Simulation experiments conducted on systems with 2, 4, and 8 GPUs on a gcloud instance with NVIDIA V100 GPUs have yielded the data, as detailed in Table 7, which is explained further in Section 4.2. It indicates that the increase in the number of GPUs, coupled with appropriate graph partitioning strategies, can lead to a reduction in total computation time required for simulation.
@@
Single GPU Challenge Solution
Traffic Atlas Simultaneous Representation of Road Network, Vehicle Speeds, and Locations inside GPU Memory (see Fig 4.1)
@@
Traffic Atlas and Vehicle States’ Markov Update GPU single instruction multiple data (SIMD) Yilmazer et al. [2014]
@@
Multi-GPU Challenge Solution
Multi-GPU Load Balancing Graph Partitioning Fig 5
Multi-GPU Communication Ghost Zone Design Fig 4
Multi-GPU Vehicle Threads Dynamic Reallocation GPU Device based Vector Fig 9
@@ 3.1 Basic Components
$l_{\text{min}}$ Minimum length of the link of the road network $v_{i}(k)$ Speed of vehicle $i$ at time step $k$
@@
$m_{i}(k)$ Probability of a mandatory lane change for vehicle $i$ at time step k $x_{i}(k)$ Distance of vehicle $i$ at time step k to an exit or intersection
$x_{0}$ Distance of a critical location to the exit or intersection $g_{\text{lead}}(k)$ Critical lead gap for a lane change at time step $k$
$g_{\text{lag}}(k)$ Critical lag gap for a lane change at time step $k$ $g_{a}$ Desired lead gap for a lane change
@@
$T$ Desired time headway $v_{\text{free}}$ Free Flow Speed
$f_{i}$ The function abstraction of IDM car following, lane change, and gap acceptance $v_{f}(k)$ Front vehicle’s speed at time step k
$l_{f}(k)$ The location of front vehicle at time step k
In this part, we introduce the vehicle dynamics used for microsimulation initially developed in Yedavalli et al. [2021]. There are three key components: car following model, lane change process, and gap acceptance.
We further present a theory of parallelizable simulation that aligns with the GPU’s SIMD architecture as shown in Eq. 1. This includes a detailed explanation of how memory is utilized not only to represent the road network but also to depict the occupancy and speed of the vehicles. Lastly, we consolidate our methodology into a comprehensive algorithm, outlined as Algorithm 1. Readers can refer to table 3 for notation in this part.
Vehicle Propagation: First, the dynamics of vehicles within our simulation is governed by the car following model as described in Equation (Car Following), which models vehicle acceleration to update vehicle locations at each time step. Second, the lane change process is encapsulated by Equation (Lane Change). For the vehicle, when the distance to an exit falls below a threshold distance $x_{0}$, it is triggered to make a mandatory lane change and the probability of such a change increases as the vehicle approaches the exit. Once a vehicle inquires for a lane change, it must assess the feasibility of this maneuver by assessing the gaps with both the leading and the lagging vehicles. This assessment is carried out according to Equation (Gap Acceptance). For LPSim, we adapt the IDM model as described in Albeaik et al. [2022], Iqbal et al. [2014], Choudhury et al. [2007]
$v_{i}(k+1)=f_{a,v_{0},\delta,s_{0},T,b}(v_{i}(k),v_{f}(k),l_{f}(k),l_{i}(k))$ (Car Following)
$m_{i}(k+1)=f_{x_{0}}(x_{i}(k))$ (Lane Change)
$g_{\text{lead}(k+1)}$ $=f_{g_{a},\alpha_{a},\alpha_{i},v_{a},\epsilon_{a}}(v_{i}(k),l_{i-n}(k),\ldots,l_{i+n}(k))$ (Gap Acceptance)
$g_{\text{lag}(k+1)}$ $=f_{g_{b},\alpha_{b},\alpha_{i},v_{b},\epsilon_{b}}(v_{i}(k),l_{i-n}(k),\ldots,l_{i+n}(k))$ (Gap Acceptance)
@@
$l_{i}(k+1)=f_{i}(\underbrace{l_{i}(k),v_{i}(k)}_{\text{Current Vehicle}},\underbrace{l_{i-n}(k),\ldots,l_{i+n}(k)}_{\text{Surrounding Vehicles}},\underbrace{l_{f}(k),v_{f}(k)}_{\text{Front Vehicle}})$ (1)
@@
Remark:
@@
    Network Information Representation: One byte can only be occupied by one vehicle, and we are not simulating the case that vehicles crash with each other.
@@
    Intersection Modeling: When vehicles approach, they will wait before the intersection and check the downstream road segment. If the downstream road segment is full, then the vehicles will be waiting at the current road segment.
@@
    Switch from one road to another road segment: When a vehicle transitions from one road segment to another, we employ atomic operations to ensure data integrity. Specifically, if two vehicles, A and B, attempt to move to road segment C simultaneously, atomic operations guarantee that only one vehicle’s thread successfully writes to segment C. This prevents the issue of both vehicles occupying the same memory address, which could otherwise result in vehicle data conflicts and potential loss of vehicle information.
As shown in Figure 4.1, the data stores inside GPU memory contains three components:
  • 1.
    Lane Map: The whole road network will be represented in GPU memory with Char type in C++ which can represent Numbers from 0 to 255. We use it in the following way:
    • (a)
      1 Byte in Memory = 1 Meter in Road Segments: This means each byte corresponds to one meter of road in the simulation.
    • (b)
      Value 255 = Not Occupied: Indicates that the byte is not occupied by any vehicle.
    • (c)
      Values 0-254 = Occupied by Vehicle: Represents that the byte is occupied by a vehicle, with the value indicating the vehicle’s speed in meters per second.
    Limitation: Using the Char type means that the maximum value we can represent is 255. Therefore, if a vehicle’s speed exceeds 254 meters per second, this framework cannot accurately represent it. Furthermore, since each byte corresponds to one meter, the precision of vehicle placement and road representation is limited to 1 meter.
    We use an adjacency list to map the directed connectivity of nodes, with list entries pointing to arrays that describe road segments (e.g., a 1D array for a 4-lane, 8-meter road segment will be the dimension of 1 X (4*8), resulting in dimensions of 1 X 32 as shown in Fig 4.1.) Using Char Type in C++ to indicate road status (occupied, empty, and vehicle speed, etc.), the real road segment is mapped in a 1D array lane map organizing lanes in order.
  • 2.
    Network Data Information: For each Edge from input, we extract and store the number of lanes, its index on lane map, upstream intersection, and downstream intersection information.
  • 3.
    Vehicle Data Information: For each vehicle, the route path and the first edge that it is going to traverse are from the input demand data after the routing. When the simulation is running, we keep a record and update each vehicle’s previous edge, current edge, next edge, and position on the current edge.
With three components above, when we are calculating each vehicle’s movement, we calculate vehicle locations at the lane map according to the edge’s index on the lane map and its position on the current edge to update the lane map
Figure 3: Data Movement Representation inside GPU
3.2 Architecture of single GPU-based traffic simulation
3.2.1 GPU Kernel Design for Vehicle and Traffic Management